OpenStudy (anonymous):

can someone help me where I'm going wrong? :( (1 pt) A producer finds that demand for his commodity obeys a linear demand equation , p+11x=30 where P is in dollars and X in thousands of units. If the cost equation is c(x)=8.8x^2+20x+2.5 , what price should be charged to maximize the profit? The price to maximize the profit is $ .

OpenStudy (anonymous):

i found the price: p=(30-11x)*x

OpenStudy (anonymous):

30x-11x^2- (8.8x^2+20x+2.5)

OpenStudy (anonymous):

= 10x-19.8x^2-2.5

OpenStudy (anonymous):

the i find the derivative: dp/dx and set it to zero.... i got x= .252525

OpenStudy (anonymous):

then i tried to plug into c(x) but it's wrong :((

OpenStudy (anonymous):

i got 8.05

OpenStudy (mertsj):

Isn't profit= revenue - cost?

OpenStudy (anonymous):

yeah

OpenStudy (mertsj):

I got 6.87 for the price

OpenStudy (anonymous):

how? it says its not correct..

OpenStudy (mertsj):

price = (30-11x)x

OpenStudy (anonymous):

right.

OpenStudy (mertsj):

So if x = .25, price = (30-11/4).25=6.68

OpenStudy (mertsj):

All depends on how many significant digits you carry for x

OpenStudy (anonymous):

i tried that one too

OpenStudy (anonymous):

its wrong.

OpenStudy (mertsj):

cost per unit is c(x)/x

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

If you want price per unit shouldn't you divide the price by 1000 since x is 1000 units?

OpenStudy (anonymous):

howd you get that

OpenStudy (mertsj):

and X in thousands of units. I read the problem you posted.

OpenStudy (anonymous):

yeah can you tell me where i went wrong?

OpenStudy (mertsj):

Well, I am not sure but I would try dividing the answers we got by 1000 and see if that is the price per unit.

OpenStudy (anonymous):

whered you get 1000 from?

OpenStudy (mertsj):

What does x represent?

OpenStudy (anonymous):

the price

OpenStudy (mertsj):

Read the problem. What does x represent.

OpenStudy (anonymous):

thousands in units.

OpenStudy (mertsj):

So when you got x = .25 that meant .25 thousands of units of 250 units. So when you got 6.87 for the price of x units that was for 1000 units. So each unit should be sold for 7 cents

OpenStudy (anonymous):

okay... so i got x= .25 and i times that by 1000...

OpenStudy (anonymous):

so whats the correct answer?

OpenStudy (anonymous):

im getting it wrong.

OpenStudy (mertsj):

So if x = .25, price = (30-11/4).25=6.68 for 250 units. 6.68/250=668 cents/250= 3 cents per unit. Try that. How many chances do you get?

OpenStudy (anonymous):

huh?

OpenStudy (mertsj):

I think the part you can't get a handle on is that x is in thousands of units. So if x is 1 that is 1000 units. If x = 2, that is 2000 units. If x = .25 that is 250 units.

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

i understand that part.

OpenStudy (mertsj):

What did you get for x?

OpenStudy (anonymous):

555003.

OpenStudy (mertsj):

Well that is not what you posted earlier.

OpenStudy (anonymous):

i plugged in 250 into c(x)

OpenStudy (anonymous):

i dont know.

OpenStudy (anonymous):

whats the answer then?

OpenStudy (anonymous):

isn't it .252525 *1000 = x= 252.525

OpenStudy (mertsj):

I don't know for sure. I wanted you to try 3 cents and I asked you how many chances you get at entering the correct answer and you didn't answer me.

OpenStudy (anonymous):

as many times.

OpenStudy (mertsj):

Did you try 3 cents?

OpenStudy (anonymous):

yeah its wrong.

OpenStudy (mertsj):

I guess I can't help you then. Sorry.